OpenStudy (anonymous):

identify the factor of the polynomial y^2+yb-ya-ab ~ i have no idea, pls help

OpenStudy (tamtoan):

look at y^2 and ya ..what's the common factor, then yb and ab ...what's the common factor ?

OpenStudy (cwrw238):

look to factor 2 psirs: y^2 - ya + yb - ab = y(y-a) + b(y - a) (y -a) is common so = (y + b)(y - a)

OpenStudy (cwrw238):

* 2 pairs